OpenStudy (anonymous):

There are two machines that produce aluminum cans. The newer machine can produce 6600 cans in 220 minutes. It takes the older machine 330 minutes to produce that many cans. If the two machines work together, how long will it take them to produce 6600 cans?

OpenStudy (anonymous):

The newer machine works at the rate of 6600/220 = 30 cans per min The older machine works at the rate of 6600/330=20 cans per min Then both machines work at the rate of 30+20=50 cans per min Soooooo, our equation to solve is: 50=6600/x

OpenStudy (anonymous):

50x=6600 x=132 min ----time it takes both machines working together to produce 6600 cans CK In 132 min, the newer machine produces 30*132=3960 cans In 132 min the older machine produces 20*132=2640 cans 3960+2640=6600 6600=6600

OpenStudy (anonymous):

does this help??

OpenStudy (anonymous):

@carlos5798

OpenStudy (anonymous):

yes